what is the rule when you got a domain of 0 1 2 3
and a range of 0 0.5 2 4.5

Answer by sravanik2005(14)

y=x^2/2 is the rule used here
since 0=0^2/2
0.5=1^2/2
2=2^2/2
4.5=3^2/2